# Bidirectional Task-Motion Planning Based on Hierarchical Reinforcement Learning for Strategic Confrontation

### v2

@@ Abstract
In swarm robotics, confrontation scenarios, including strategic confrontations, require efficient decision-making that integrates discrete commands and continuous actions. Traditional task and motion planning methods separate decision-making into two layers, but their unidirectional structure fails to capture the interdependence between these layers, limiting adaptability in dynamic environments. Here, we propose a novel bidirectional approach based on hierarchical reinforcement learning, enabling dynamic interaction between the layers. This method effectively maps commands to task allocation and actions to path planning, while leveraging cross-training techniques to enhance learning across the hierarchical framework. Furthermore, we introduce a trajectory prediction model that bridges abstract task representations with actionable planning goals. In our experiments, it achieves over 80% in confrontation win rate and under 0.01 seconds in decision time, outperforming existing approaches. Demonstrations through large-scale tests and real-world robot experiments further emphasize the generalization capabilities and practical applicability of our method.
@@ I INTRODUCTION
Recent advances in artificial intelligence lead to significant progress in robotics [1, 2], with particular attention given to robotic swarm confrontations [3, 4]. In these scenarios, AI-driven decision-making empowers robots to engage in strategic tasks such as pursuit-evasion [5], defense-attack [6], and strategic confrontations [7]. These applications highlight the need for effective decision-making mechanisms capable of handling a hybrid process, including discrete commands and continuous actions [8], which are central to the success of modern robotic systems.
@@
Traditional methods for solving such problems rely heavily on task and motion planning (TAMP) [8, 9, 10, 11, 12, 13], which divides the problem into two distinct decision-making layers. Although conventional TAMP methods [8, 9] demonstrate effectiveness by organizing tasks within detailed planning frameworks, they suffer limitations due to their reliance on expert knowledge [10]. The complexity of modeling and task specification renders these methods less accessible for non-expert users. Heuristic algorithms [11, 12] attempt to address these challenges by adaptively searching for feasible solutions, but the independent nature of the hierarchical architecture prevents the upper layer from verifying the enforceability of commands transferred to the lower layer [13]. Furthermore, TAMP methods struggle with generalization, as evidenced by the frequent need for repeated adjustments [14] when adapting to dynamic environments or scaling to larger robot populations.
Figure 1: Simulations and real-robot systems in a confrontation scenario.
@@
In contrast, deep reinforcement learning (DRL) [15, 16] emerges as a promising alternative by using a single-layer framework [17] that learns optimal decision-making strategies through end-to-end training. It aims to maximize cumulative rewards, fine-tuning its decision network to produce effective strategies. However, in complex environments [18], DRL encounters challenges due to its focus on global objectives, leading to sparse-reward scenarios [19]. This results in practical limitations when applied to intricate real-world tasks.
@@
To address these limitations, hierarchical reinforcement learning (HRL) [20] provides a more flexible solution by decomposing the complex task into two independently trainable networks, similar to the TAMP approach. It fosters a dynamic feedback loop between the two layers, enabling them to communicate and refine their strategies as they learn. In HRL, the upper layer divides the task into temporal segments and assigns tailored rewards to guide the lower layer. The method allows the upper layer to refine its task decomposition based on the learning progress of the lower layer, offering a more adaptable framework to tackle complex problems [21, 22]. This hierarchical structure enables more detailed problem-solving, particularly in scenarios where discrete commands and continuous actions are closely intertwined.
@@
In this paper, we propose an innovative HRL method to tackle the unique challenges of swarm confrontation. Unlike previous approaches, our method introduces a two-layer interaction mechanism that enhances both the accuracy and stability of decision-making. Specifically, we integrate task allocation (for discrete commands) and path planning (for continuous actions) within a bidirectional closed-loop architecture. This structure enables effective communication between the two layers, ensuring the enforceability of assigned tasks. Additionally, we introduce a trajectory prediction model that bridges the gap between abstract task representations and executable planning targets, such as predicting the most likely location of enemies in an evasion task. Our experimental results demonstrate that our approach outperforms baseline methods, including traditional TAMP approaches and single-layer DRL, especially in terms of generalization across different scales. Through ablation studies, we highlight the critical importance of the interaction mechanism and show that our method exhibits superior adaptability in real-robot systems illustrated in Fig. 1.
@@
We organize the rest of the paper as follows. Section II outlines the preliminaries of goal-conditioned HRL and formulates the strategic confrontation problem. Section III elaborates on the proposed HRL method and its implementation. Section IV presents the experimental evaluations of our approach. Finally, Section V concludes the paper.
@@ II Preliminaries and Problem Description
II-A Goal-conditioned HRL
Let $\mathbb{R}$ denote the set of real numbers, $\mathbb{R}^{+}$ denote the set of positive real numbers, and $\mathbb{Z}$ denote the set of integers. We define a long-horizon planning task for $N$ agents using a goal-conditioned Markov Decision Process (goal-conditioned MDP), represented as an eight-tuple $\left(N,\mathbb{S},\mathbb{G},\mathbb{O},\mathbb{A},P,R,\gamma\right)$, where:
@@ II-A Goal-conditioned HRL
  • $\boldsymbol{z}_{i}\in\mathbb{O}$ represents the observation of the $i$-th agent and $\boldsymbol{z}=\left(\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{N}\right)$ denotes the joint observations.
  • $\boldsymbol{a}_{i}\in\mathbb{A}$ represents the action of the $i$-th agent and $\boldsymbol{a}=\left(\boldsymbol{a}_{1},\cdots,\boldsymbol{a}_{N}\right)$ denotes the joint actions.
@@
The subgoal $\boldsymbol{q}$ and the goal $\boldsymbol{g}$ are defined in the goal space $\mathbb{G}$, which is considered a subspace of $\mathbb{S}$, with a mapping function $\mathcal{\varphi}:\mathbb{S}\rightarrow\mathbb{G}$. HRL typically involves two-layer policies: an upper-layer policy $\pi_{u}\left(\boldsymbol{q}\left(t\right)\left|\boldsymbol{z}\left(t\right),\boldsymbol{g}\right.\right)$ and a lower-layer policy $\pi_{l}\left(\boldsymbol{a}\left(t\right)\left|\boldsymbol{z}\left(t\right),\boldsymbol{q}\left(t\right)\right.\right)$. We assume that each subtask consists of $h$ steps. The upper layer predicts the subgoal $\boldsymbol{q}_{m},m=0,1,\cdots,T_{h}$, where $T_{h}=T/h$ and $T$ is the total time. Upon receiving the observation $\boldsymbol{z}_{m+n}$, the lower layer attempts to reach the subgoal by executing the appropriate action $\boldsymbol{a}_{n}$. Starting from the initial state $\boldsymbol{s}_{m}$, the state-action trajectory for the $m$-th subtask can be represented as $\tau_{m}=\{\left(\boldsymbol{s}_{m+n},\boldsymbol{a}_{n},\boldsymbol{s}_{m+n+1}\right)\left|\boldsymbol{s}_{m},\boldsymbol{q}_{m}\right.\}^{h-1}_{n=0}$.
To ensure the enforceability of the predicted subgoal, the upper-layer policy is guided by the reward feedback from the environment. The reward function and optimization objective in the upper layer can be described as follows:
@@
II-B Definition of strategic confrontation
This study models strategic confrontation as a dynamic process that involves task allocation and path planning in a static-obstacle environment. In this scenario, the blue and red agents compete against each other, with the one that eliminates all opponents being declared the winner. To simplify the agents’ movement, we consider the confrontation in two-dimensional space. Let $\boldsymbol{p}$ and $\boldsymbol{v}$ denote the position and velocity vectors in two dimensions, respectively. Agents on both sides are subject to the same constraints on their abilities. For the $i$-th agent $\boldsymbol{u}_{i}=\left(\boldsymbol{p}_{i},\boldsymbol{v}_{i},ta_{i}\right)$, we define $\mathbb{U}_{a}=\left\{u_{a,j}|j=1,...,N_{a}\right\}$ and $\mathbb{U}_{e}=\left\{u_{e,k}|k=1,...,N_{e}\right\}$ as the sets of $N_{a}$ allies and $N_{e}$ enemies, respectively. Here, $ta$ represents the task that is being performed at the current moment, which can be searching $\left(ta=-3\right)$, escaping $\left(ta=-2\right)$, supporting $\left(ta=-1\right)$, or chasing $\left(ta=0\right)$. At each time step, every agent first assigns a task and then performs path planning based on the task goal to ensure obstacle avoidance.
As shown in Fig. 1, each agent has an attack range $D_{1}$ modeled as a sector with radius $\rho_{1}$ and angle $\theta_{1}$. When an opponent enters the attack range $D_{1}$, the agent automatically launches an attack. At this point, the probability of the bullet hitting and causing the opponent to die is $\epsilon$. By defining an avoidance range $D_{2}$ with radius $\rho_{2}$, the agent plans feasible paths to avoid collisions with obstacles. In this work, the $i$-th agent gathers information about neighboring allies and enemies through communication and perception, respectively. We define the communication and perception ranges as a rectangle with length $d_{1}$ and width $d_{2}$. While obstacles obstruct perception, they do not interfere with communication. When the $j$-th agent is within the communication and perception ranges of the $i$-th agent, we label it as $j\in\mathbb{N}_{i}$, where $\mathbb{N}_{i}$ denotes the set of agents observable by the $i$-th agent.
Figure 2: Overview of the proposed method.
@@ III METHODOLOGY
In this section, we present the HRL method to address the strategic confrontation problem in a static-obstacle environment. The lower layer utilizes a multi-agent deep deterministic policy gradient (MADDPG) algorithm for path planning based on subgoals. The upper layer employs a decentralized deep Q-learning (DQN) algorithm for task allocation. This method integrates the cumulative rewards from the lower layer into the upper layer, establishing an interaction mechanism between the two layers. In addition, we employ a trajectory prediction model to bridge abstract task representations with concrete planning goals. Finally, we introduce a cross-training strategy to enhance the training efficiency and stability of the proposed method. An overview of our method is shown in Fig. 2.
@@ III-A Lower Layer for Path Planning
$\begin{array}[]{c}\boldsymbol{z}_{l,i}=\left(\boldsymbol{p}_{i},\boldsymbol{p}_{j},\boldsymbol{v}_{i},\boldsymbol{v}_{j},\boldsymbol{p}_{o,c}\right),\end{array}$ (5)
where $j,c\in\mathbb{N}_{i}$. The action consists of the velocity magnitude $\lVert\boldsymbol{v}\rVert\in[0,\lVert\boldsymbol{v}\rVert_{\max}]$ and velocity direction $\psi\in[-\pi,\pi]$ of the $i$-th agent, denoted as
@@
[…]
@@
We design an avoidance reward to guide agents in avoiding collisions with obstacles and other agents. Specifically, if the $i$-th agent collides ($\lVert\boldsymbol{p}_{i}-\boldsymbol{p}_{j}\rVert<2\rho_{2}$ or $\lVert\boldsymbol{p}_{i}-\boldsymbol{p}_{o,c}\rVert<\rho_{2}+\rho_{3}$), the avoidance reward $r_{a}=-1$. Otherwise, $r_{a}=0$. This reward, therefore, actively promotes collision avoidance and safe navigation. Although avoiding collisions is crucial, agent pathfinding must also prioritize minimizing path length to ensure efficiency. The upper layer introduces an intrinsic reward $r_{b}$, which is defined as follows:
$\displaystyle r_{b}=-\frac{\lVert\boldsymbol{q}_{i}\left(\left\lfloor t/h\right\rfloor\right)-\boldsymbol{p}_{i}\left(t\right)\rVert}{\lVert\boldsymbol{q}_{i}\left(\left\lfloor t/h\right\rfloor\right)-\boldsymbol{p}_{i}\left(\left\lfloor t/h\right\rfloor\right)\rVert}.$ (8)
@@
$R_{l}\left(\boldsymbol{s}_{m+n},\boldsymbol{a}_{n},\boldsymbol{q}_{m}\right)=\epsilon_{1}r_{a}+\left(1-\epsilon_{1}\right)r_{b}.$ (9)
where $\epsilon_{1}$ is the weighting factor that balances avoidance reward and intrinsic reward. It encourages the agent to navigate safely while minimizing the path length, facilitating efficient completion of its target-reaching task.
@@ III-B Upper Layer for Task Allocation
The goal $\boldsymbol{g}$ of the confrontation is to eliminate as many enemies as possible. Given the limited observations in the confrontation scenario, the local observation of the $i$-th agent includes information about observable allies and enemies, expressed as
@@
where $j,k\in\mathbb{N}_{i}$. Furthermore, the tasks of enemies $\boldsymbol{u}_{e,k}$ are unknown to the $i$-th agent.
Based on the goal and observation, the agent chooses to either support an ally that is escaping or chasing, or to chase an enemy at a relative disadvantage. The upper layer generates the subgoal for the $i$-th agent by determining the target position defined as follows:
@@
where $ta_{a,j}=-2$ or $ta_{a,j}=-1$, and $Adv\left(\boldsymbol{u}_{i},\boldsymbol{u}_{e,k}\right)>0$. The advantage function $Adv$ depends on the positions and velocities of $\boldsymbol{u}_{i}$ and $\boldsymbol{u}_{e,k}$ as follows:
$Adv\left(\boldsymbol{u}_{i},\boldsymbol{u}_{e,k}\right)=\cos\left(\boldsymbol{v}_{i},\boldsymbol{p}_{e,k}-\boldsymbol{p}_{i}\right),$ (12)
where $\cos$ denotes the cosine function. The selection of enemies and allies determines whether the agent is engaged in a chasing or supporting task. If neither allies nor enemies, as described above, are within the communication and perception ranges, the agent will perform a searching or escaping task and invoke the trajectory prediction model. After the lower layer executes $h$ steps, we feed the cumulative rewards to the upper layer, which adjusts its strategy based on the performance of the lower layer through bidirectional feedback. Therefore, the reward function of the upper layer in (1) is improved as

### v3

# Bidirectional Task–Motion Planning Based on Hierarchical Reinforcement Learning for Strategic Confrontation

@@ Abstract
In swarm robotics, confrontation scenarios, including strategic confrontations, require efficient decision–making that integrates discrete commands and continuous actions. Traditional task and motion planning methods separate decision–making into two layers, but their unidirectional structure fails to capture the interdependence between these layers, limiting adaptability in dynamic environments. Here, we propose a novel bidirectional approach based on hierarchical reinforcement learning, enabling dynamic interaction between the layers. This method effectively maps commands to task allocation and actions to path planning, while leveraging cross–training techniques to enhance learning across the hierarchical framework. Furthermore, we introduce a trajectory prediction model that bridges abstract task representations with actionable planning goals. In our experiments, it achieves over 80% in confrontation win rate and under 0.01 seconds in decision time, outperforming existing approaches. Demonstrations through large–scale tests and real–world robot experiments further emphasize the generalization capabilities and practical applicability of our method.
@@ I INTRODUCTION
Recent advances in artificial intelligence lead to significant progress in robotics [1, 2], with particular attention given to robotic swarm confrontations [3, 4]. In these scenarios, AI–driven decision–making empowers robots to engage in strategic tasks such as pursuit–evasion [5], defense–attack [6], and strategic confrontations [7]. These applications highlight the need for effective decision–making mechanisms capable of handling a hybrid process, including discrete commands and continuous actions [8], which are central to the success of modern robotic systems.
@@
Traditional methods for solving such problems rely heavily on task and motion planning (TAMP) [9], which divides the problem into two distinct decision–making layers. Although conventional TAMP methods [8, 10] demonstrate effectiveness by organizing tasks within detailed planning frameworks, they suffer limitations due to their reliance on expert knowledge . The complexity of modeling and task specification renders these methods less accessible for non–expert users. Heuristic algorithms [11, 12] attempt to address these challenges by adaptively searching for feasible solutions, but the independent nature of the hierarchical architecture prevents the upper layer from verifying the enforceability of commands transferred to the lower layer. Furthermore, TAMP methods struggle with generalization, as evidenced by the frequent need for repeated adjustments [13, 14] when adapting to dynamic environments or scaling to larger robot populations.
Figure 1: Simulations and real–robot systems in a confrontation scenario.
@@
In contrast, deep reinforcement learning (DRL) [15, 16] emerges as a promising alternative by using a single–layer framework [17] that learns optimal decision–making strategies through end–to–end training. It aims to maximize cumulative rewards, fine–tuning its decision network to produce effective strategies. However, in complex environments [18], DRL encounters challenges due to its focus on global objectives, leading to sparse–reward scenarios [19]. This results in practical limitations when applied to intricate real–world tasks.
@@
To address these limitations, hierarchical reinforcement learning (HRL) [20] provides a more flexible solution by decomposing the complex task into two independently trainable networks, similar to the TAMP approach. It fosters a dynamic feedback loop between the two layers, enabling them to communicate and refine their strategies as they learn. In HRL, the upper layer divides the task into temporal segments and assigns tailored rewards to guide the lower layer. The method allows the upper layer to refine its task decomposition based on the learning progress of the lower layer, offering a more adaptable framework to tackle complex problems [21, 22]. This hierarchical structure enables more detailed problem–solving, particularly in scenarios where discrete commands and continuous actions are closely intertwined.
@@
In this paper, we propose an innovative HRL method to tackle the unique challenges of swarm confrontation. Unlike previous approaches, our method introduces a two–layer interaction mechanism that enhances both the accuracy and stability of decision–making. Specifically, we integrate task allocation (for discrete commands) and path planning (for continuous actions) within a bidirectional closed–loop architecture. This structure enables effective communication between the two layers, ensuring the enforceability of assigned tasks. Additionally, we introduce a trajectory prediction model that bridges the gap between abstract task representations and executable planning targets, such as predicting the most likely location of enemies in an evasion task. Our experimental results demonstrate that our approach outperforms baseline methods, including traditional TAMP approaches and single–layer DRL, especially in terms of generalization across different scales. Through ablation studies, we highlight the critical importance of the interaction mechanism and show that our method exhibits superior adaptability in real–robot systems illustrated in Fig. 1.
@@
We organize the rest of the paper as follows. Section II outlines the preliminaries of goal–conditioned HRL and formulates the strategic confrontation problem. Section III elaborates on the proposed HRL method and its implementation. Section IV presents the experimental evaluations of our approach. Finally, Section V concludes the paper.
@@ II Preliminaries and Problem Description
II-A Definition of strategic confrontation
This study models strategic confrontation as a dynamic process that involves task allocation and path planning in a static–obstacle environment. In this scenario, the blue and red agents compete against each other, with the one that eliminates all opponents being declared the winner. To simplify the agents’ movement, we consider the confrontation in two–dimensional space. Let $\boldsymbol{p}$ and $\boldsymbol{v}$ denote the position and velocity vectors in two dimensions, respectively. Agents on both sides are subject to the same constraints on their abilities. For the ii–th agent $\boldsymbol{u}_{i}=\left(\boldsymbol{p}_{i},\boldsymbol{v}_{i},f_{i}\right)$, we define $\mathbb{U}_{a}=\left\{u_{a,j}|j=1,...,N_{a}\right\}$ and $\mathbb{U}_{e}=\left\{u_{e,k}|k=1,...,N_{e}\right\}$ as the sets of $N_{a}$ allies and $N_{e}$ enemies, respectively. Here, $f_{i}$ represents the task currently performed by the ii–th agent, defined as searching ($f_{i}=-3$), escaping ($f_{i}=-2$), supporting ($f_{i}=-1$), or chasing ($f_{i}=0$). At each time step, every agent assigns a task and then performs path planning based on the task goal to ensure obstacle avoidance.
As shown in Fig. 1, each agent has an attack range $D_{1}$ modeled as a sector with radius $\rho_{1}$ and angle $\theta_{1}$. When an opponent enters the attack range $D_{1}$, the agent automatically launches an attack. At this point, the probability of the bullet hitting and causing the opponent to die is $\epsilon$. By defining an avoidance range $D_{2}$ with radius $\rho_{2}$, the agent plans feasible paths to avoid collisions with obstacles. In this work, the ii–th agent gathers information about neighboring allies and enemies through communication and perception, respectively. We define the communication and perception ranges as a rectangle with length $d_{1}$ and width $d_{2}$. While obstacles obstruct perception, they do not interfere with communication. When the jj–th agent is within the communication and perception ranges of the ii–th agent, we label it as $j\in\mathbb{N}_{i}$, where $\mathbb{N}_{i}$ denotes the set of agents observable by the ii–th agent. This confrontation setting provides the foundational context for the goal–conditioned HRL framework.
Figure 2: Overview of the proposed framework integrating task allocation (left), trajectory prediction (top right), and path planning (bottom right) modules.
II-B Goal–conditioned HRL
Let $\mathbb{R}$ denote the set of real numbers, $\mathbb{R}^{+}$ the set of positive real numbers, and $\mathbb{Z}$ the set of integers. We formulate a long–horizon planning task for $N$ agents as a goal–conditioned Markov Decision Process (goal–conditioned MDP), represented by the eight–tuple $\left(N,\mathbb{S},\mathbb{G},\mathbb{O},\mathbb{A},P,R,\gamma\right)$, where:
@@ II-B Goal–conditioned HRL
  • $\boldsymbol{z}_{i}\in\mathbb{O}$ represents the observation of the ii–th agent and $\boldsymbol{z}=\left(\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{N}\right)$ denotes the joint observations.
  • $\boldsymbol{a}_{i}\in\mathbb{A}$ represents the action of the ii–th agent and $\boldsymbol{a}=\left(\boldsymbol{a}_{1},\cdots,\boldsymbol{a}_{N}\right)$ denotes the joint actions.
@@
[…]
To ensure the enforceability of the predicted subgoal, the upper–layer policy is guided by the reward feedback from the environment. The reward function and optimization objective in the upper layer can be described as follows:
@@ III METHODOLOGY
In this section, we present the HRL method to address the strategic confrontation problem in a static–obstacle environment. An overview of our method is shown in Fig. 2. The lower layer utilizes a multi–agent deep deterministic policy gradient (MADDPG) algorithm [5] for path planning based on subgoals. The upper layer employs a decentralized deep Q–learning (DQN) algorithm [7] for task allocation. This method integrates the cumulative rewards from the lower layer into the upper layer, establishing an interaction mechanism between the two layers. In addition, we employ a trajectory prediction model to bridge abstract task representations with concrete planning goals. Finally, we introduce a cross–training strategy to enhance the training efficiency and stability of the proposed method.
@@ III-A Lower Layer for Path Planning
$\begin{array}[]{c}\boldsymbol{z}_{l,i}=\left(\boldsymbol{p}_{i},\boldsymbol{p}_{j},\boldsymbol{v}_{i},\{\boldsymbol{v}_{j}\}_{j\in\mathbb{N}_{i}},\{\boldsymbol{p}_{o,c}\}_{c\in\mathbb{N}_{i}}\right).\end{array}$ (5)
The action consists of the velocity magnitude $\lVert\boldsymbol{v}\rVert\in[0,\lVert\boldsymbol{v}\rVert_{\max}]$ and velocity direction $\psi\in[-\pi,\pi]$ of the ii–th agent, denoted as
@@
where $\lVert\cdot\rVert$ denotes the Euclidean norm and $\lVert\boldsymbol{v}\rVert_{\max}$ denotes the maximum velocity magnitude. Based on the current state and actions, we can calculate the position at the next time step using the first–order dynamics model
@@
We design an avoidance reward to guide agents in avoiding collisions with obstacles and other agents. Specifically, if the ii–th agent collides ($\lVert\boldsymbol{p}_{i}-\boldsymbol{p}_{j}\rVert<2\rho_{2}$ or $\lVert\boldsymbol{p}_{i}-\boldsymbol{p}_{o,c}\rVert<\rho_{2}+\rho_{3}$), the avoidance reward $r_{i,a}=-1$. Otherwise, $r_{i,a}=0$. This reward, therefore, actively promotes collision avoidance and safe navigation. Although avoiding collisions is crucial, agent pathfinding must also prioritize minimizing path length to ensure efficiency. The upper layer introduces an intrinsic reward $r_{i,b}$, which is defined as follows:
$\displaystyle r_{i,b}=-\frac{\lVert\boldsymbol{q}_{i}\left(\left\lfloor t/h\right\rfloor\right)-\boldsymbol{p}_{i}\left(t\right)\rVert}{\lVert\boldsymbol{q}_{i}\left(\left\lfloor t/h\right\rfloor\right)-\boldsymbol{p}_{i}\left(\left\lfloor t/h\right\rfloor\right)\rVert}.$ (8)
@@
$R_{l}\left(\boldsymbol{s}_{m+n},\boldsymbol{a}_{n},\boldsymbol{q}_{m}\right)=\sum_{i}^{N}(\epsilon_{1}r_{i,a}+\left(1-\epsilon_{1}\right)r_{i,b}).$ (9)
where $\epsilon_{1}$ is the weighting factor that balances avoidance reward and intrinsic reward. It encourages the agent to navigate safely while minimizing the path length, facilitating efficient completion of its target–reaching task.
@@ III-B Upper Layer for Task Allocation
The goal $\boldsymbol{g}$ of the confrontation is to eliminate as many enemies as possible. Given the limited observations in the confrontation scenario, the local observation of the ii–th agent includes information about observable allies and enemies, expressed as
@@
where $j,k\in\mathbb{N}_{i}$. Furthermore, the tasks of enemies $\boldsymbol{u}_{e,k}$ are unknown to the ii–th agent.
Based on the goal and observation, the agent chooses to either support an ally that is escaping or chasing, or to chase an enemy at a relative disadvantage. The upper layer generates the subgoal for the ii–th agent by determining the target position defined as follows:
@@
where $f_{a,j}=-2$ or $f_{a,j}=0$, and $Adv\left(\boldsymbol{u}_{i},\boldsymbol{u}_{e,k}\right)>0$. The advantage function $Adv$ depends on the positions and velocities of $\boldsymbol{u}_{i}$ and $\boldsymbol{u}_{e,k}$, denoted as $Adv\left(\boldsymbol{u}_{i},\boldsymbol{u}_{e,k}\right)=\cos\theta_{i,k}$, where $\theta_{i,k}$ represents the angle between the vectors $\boldsymbol{v}_{i}$ and $\boldsymbol{p}_{e,k}-\boldsymbol{p}_{i}$. The selection of enemies and allies determines whether the agent is engaged in a chasing or supporting task. If neither allies nor enemies, as described above, are within the communication and perception ranges, the agent will perform a searching or escaping task and invoke the trajectory prediction model. After the lower layer executes $h$ steps, we feed the cumulative rewards to the upper layer, which adjusts its strategy based on the performance of the lower layer through bidirectional feedback. Therefore, the reward function of the upper layer in (1) is improved as